What is the perimeter and area of the rectangle the top side is 3+6 times the square root of 2 yards and the right side is the square root of 8 yards.
Answer by mananth(16949)

What is the perimeter and area of the rectangle the top side is 3+6 times the square root of 2 yards and the right side is the square root of 8 yards.
length = 3+6sqrt2
width= sqrt8
perimeter = 2(3+6sqrt2) + 2 (sqrt8)
= 6+12sqrt2+2sqrt8
=6+12sqrt2+2*2 sqrt2----------------- 2sqrt8 = 2sqrt (2*4)=2sqrt2*sqrt4 =4sqrt2
=6+12sqrt2+4sqrt2
=6+16sqrt2 yards = perimeter
Area = L*W
= (3+6sqrt2)*2sqrt8
=(3+6sqrt2)*4sqrt2
=4sqrt2 +24sqrt2
=28sqrt2 square yards is the area